OBSERVATION OF TWO COMBINATION BANDS INVOLVING TORSION AND ASYMMETRIC BEND-ING MODES OF THE NON-POLAR N_{2}O DIMER

\underline{M. DEHGHANY}, M. AFSHARI, N. MOAZZEN-AHMADI, Department of Physics and Astronomy, 2500 University Drive NW, Calgary, ALberta T2N 1N4, Canada; A.R.W. MCKELLAR, Steacie Institute for Molecular Sciences, National Research Council of Canada, Ottawa, Ontario K1A OR6, Canada.

The non-polar N_{2}O dimer has four low frequency intermolecular vibrational modes: van der Waals stretch (A_{g} symmetry), symmetric and asymmetric in-plane bends (A_{g} and B_{u} symmetry, respectively), and out-of-plane torsion (A_{u} symmetry). There are two possible intramolecular dimer vibrations which correlate with the ν_{1} monomer stretch. One of these is an in-phase vibration of the two monomers, which has A_{g} symmetry and is not infrared active. The other is an out-of-phase vibration, which has B_{u} symmetry and gives rise to the infrared active dimer band at 2229.48 cm^{−1}. In the present work, we report the observation of an N_{2}O dimer band with c-type rotational structure, assigned as a combination of the A_{g} intramolecular N_{2}O ν_{1} stretching vibration and the A_{u} intermolecular out-of-plane dimer torsional vibration. The vibrational origin for this band is measured to be 2249.360 cm^{−1}, giving an estimated torsional frequency of 21.5 cm^{−1}. A second combination band with a- and b-type rotational transitions has also been measured and assigned as the combination of the same A_{g} intramolecular vibration and the B_{u} asymmetric van der Waals bend. The measured band center is 2264.37 cm^{−1}, giving an asymmetric bending frequency of about 36.5 cm^{−1}. A search for the other two combination bands is currently underway.